JosephGunnels:

Given: f(x) = x - 7 and h(x) = 2x + 3 Write the rule for f(h(x)). f(h(x)) = 2x - 7 f(h(x)) = 2x - 4 f(h(x)) = 3x - 4 f(h(x)) = 3x - 7

JosephGunnels:

I need help

XioGonz:

f(h(x)) = 2x - 4

JosephGunnels:

thank you

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JosephGunnels thank you \(\color{#0cbb34}{\text{End of Quote}}\) Np and How I got it f(h(x)) = (2x + 3) - 7 = 2x + 3 - 7 = 2x - 4

JosephGunnels:

ok

supie:

Ok, so the functions that were given were \[f(x)=x+7\] \[h(x)=2x+3\] The question is asking us to find the value of \[f(h(x))\] \[f(h(x))=f(2x+3)\] \[f(x)=x-7\] Substitute x=(2x+3) in function f(x).... \[f(2x+3)=(2x+3)-7\] \[f(2x+3)=2x+3=7\] \[f(2x+3)=2x-4\] so now we must substitute the value of f(2x+3) in equation (1). \[f(h(x))=2x-4\]